OpenStudy (anonymous):

What is 2 (square root) 45 expressed in simplest radical form?

OpenStudy (lgbasallote):

\[2 \sqrt{45} = 2 \times \sqrt 9 \times \sqrt 5\]

OpenStudy (lgbasallote):

wait let me do that step-by-step... \[2\sqrt{45} = 2\sqrt{9 \times 5} = 2 \times \sqrt 9 \times \sqrt 5\]

OpenStudy (lgbasallote):

so what is \(\sqrt 9?\)

OpenStudy (anonymous):

3?

OpenStudy (lgbasallote):

yup so that is \[2 \times 3 \times \sqrt 5\] what's 2 x 3?

OpenStudy (anonymous):

6

OpenStudy (lgbasallote):

so the simplified form is \(6\sqrt{5}\) got that?

OpenStudy (anonymous):

yes . thank you

OpenStudy (lgbasallote):

great! ^_^